# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ IV-B Security Vulnerability of CDKS
[…]
    =\ = $\sum_{i=1}^{n}(c_{0}^{i}+c_{1}^{i}\cdot s_{i})+\sum_{i=1}^{n}e_{i}\approx\sum_{i=1}^{n}\mu_{i}\ (\text{mod}\ Q)$
@@
Although the above process ensures correct parameter aggregation, it exposes local client parameters even under the honest-but-curious assumption, where both the server and the clients are assumed to faithfully execute the FL protocol while being curious to infer the privacy data of other clients based on the information it observes during the execution of the protocol. This security vulnerability stems from the ciphertext expansion and distributed decryption processes. Specifically, each component $\bar{c}_{i}$ ($i\in[1,n]$) of the expanded ciphertext $\overline{ct}=(\bar{c}_{0},\bar{c}_{1},\bar{c}_{2},\cdots,\bar{c}_{n})$ is associated with only the fresh ciphertext $c_{1}^{i}$ of client $i$. This design facilitates decryption correctness: during partial decryption, each client $i$ simply uses its secret key $s_{i}$ to decrypts its corresponding component $\bar{c}_{i}$, obtaining the partial decryption result $\nu_{i}=c_{1}^{i}\cdot s_{i}+e_{i}\pmod{Q}$. In the full decryption phase, the server (or all clients) merges the shared values as $\bar{c}_{0}+\sum_{i=1}^{n}\nu_{i}\approx\sum_{i=1}^{n}\mu_{i}\pmod{Q}$, since each individual decryption step satisfies $c_{0}^{i}+\nu_{i}\approx\mu_{i}$, and the first component $\bar{c}_{0}=\sum_{i=1}^{n}c_{0}^{i}$. Thus, the correctness of the full decryption is ensured. However, this simple mechanism, while guaranteeing correct decryption, introduces a critical security vulnerability. Since $c_{0}^{i}$, as part of the fresh ciphertext $ct_{i}=(c_{0}^{i},c_{1}^{i})$, is public, and the partial decryption result $\nu_{i}$ is shared to either the server or the clients for full decryption, any participant with access to $\nu_{i}$ can directly recover the client $i$’s plaintext via $c_{0}^{i}+\nu_{i}\approx\mu_{i}\pmod{Q}$.
@@
V New Multi-Key Variants of CKKS and BFV
To achieve our secure multi-key homomorphic encryption (SMHE), we retain the foundational structure of CDKS but introduce a masking scheme. This masking scheme ensures both the security and correctness of homomorphic addition, thereby enabling secure aggregation applications such as FL. Furthermore, ciphertexts processed through homomorphic addition with the masking scheme can still be directly evaluated by the homomorphic multiplication operations of CDKS without additional requirements or auxiliary computations.
V-A A Masking Scheme for CKKS and BFV
We design a masking scheme tailored for CKKS and BFV, highlighting its role as a critical component in MKHE framework. Essentially, a masking scheme enables the use of a CKKS/BFV public key $pk$ (with an associated secret key $sk=(1,s)$) and a plaintext $\mu$ to generate a triple $(ct,cz,\bm{\Gamma})$. Here, $ct=(c_{0},c_{1})$ represents CKKS/BFV encryption of $\mu$ under $pk$, while $cz=(z_{0},z_{1})$ and $\bm{\Gamma}=(\bm{\varsigma}_{0},\bm{\varsigma}_{1})$ act as auxiliary information with two key properties: (1) the triple $(ct,cz,\bm{\Gamma})$ provides computational hiding for $\mu$, similar to $ct$ alone, and (2) when provided with another CKKS/BFV public key $pk^{\prime}$ (corresponding to a secret key $sk^{\prime}=(1,s^{\prime})$), it becomes feasible to construct a pair of polynomials $cx=(x_{0},x_{1})\in R_{Q}^{2}$ according to $\bm{\Gamma}$, such that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ and $\langle sk,ct\rangle+\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx\mu$. The latter also implies that $\langle(1,s,s^{\prime}),(c_{0}+x_{0}+z_{0},c_{1}+x_{1},z_{1})\rangle=\langle\overline{sk},\overline{ct}\rangle\approx\mu$. Here, $\overline{sk}=(1,s,s^{\prime})$ is the decryption key and $\overline{ct}=(c_{0}+x_{0}+z_{0},c_{1}+x_{1},z_{1})$ is a masked ciphertext.
@@ V-A A Masking Scheme for CKKS and BFV
$(pp,pk,\texttt{UniEnc}(\mu,pk))$ $\stackrel{{\scriptstyle\text{comp}}}{{\approx}}(pp,pk,\texttt{UniEnc}(\mu^{\prime},pk)),$
$(pp,pk,\texttt{MaskEnc}(r,pk))$ $\stackrel{{\scriptstyle\text{comp}}}{{\approx}}(pp,pk,\texttt{MaskEnc}(r^{\prime},pk)),$
[…]
@@
[…]
    ${\color[rgb]{0,0,1}=_{\scriptsize\text{CKKS}}w\cdot pk+(\mu+e_{w_{0}},e_{w_{1}})\pmod{Q}.}$
    ${\color[rgb]{1,0,0}=_{\scriptsize\text{BFV}}w\cdot pk+(\lfloor(Q/t)\cdot\mu\rceil+e_{w_{0}},e_{w_{1}})\pmod{Q}.}$
@@
      Sample $e_{r_{0}},e_{r_{1}}\leftarrow D_{\sigma}$.
@@
      Encrypt the value 00 with $r$ and $pk$ using CKKS/BFV encryption algorithm and output
      $\begin{split}cz&\leftarrow\texttt{FHE.Encrypt}(pk,0;r)\\ &=r\cdot pk+(e_{r_{0}},e_{r_{1}})\pmod{Q}.\\ \end{split}$ (6)
    3. 3.
@@
The $\texttt{Extend}^{*}$ algorithm simplifies the process of running the Extend algorithm for all $n$ public key $\{pk_{i}\}_{i\in[1,n]}$. Rather than executing $\texttt{Extend}(\bm{\Gamma},pk,pk_{1})+\cdots+\texttt{Extend}(\bm{\Gamma},pk,pk_{n})=\sum_{i\in[1,n]}(b_{i}-b)\boxdot\bm{\Gamma}$, it consolidates these computations into a single execution of $\texttt{Extend}^{*}(\bm{\Gamma},pk,\{pk_{i}\}_{i\in[1,n]})$.
Semantic Security. The attacker’s view consists of the distribution $\left(pp,pk,ct,cz,\bm{\Gamma}\right)$, where $pp\leftarrow\texttt{SMHE.Setup}(1^{\lambda})$, $(sk,pk)\leftarrow\texttt{SMHE.}\texttt{Keygen}(pp)$, $ct\leftarrow\texttt{UniEnc}(\mu,pk)$, and $(cz,\bm{\Gamma})\leftarrow\texttt{MaskEnc}(r,pk)$, where $\mu\in R_{Q}$.
The semantic security proof leverages the security of CKKS/BFV encryption and follows these steps: (1) Modify $\bm{\Gamma}$ to the gadget encryption of 00 instead of the gadget encryption of $r$. This is justified by the semantic security of CKKS/BFV encryption. (2) Replace $ct$ and $cz$ with encryptions of random messages. Similarly, this step relies on the security of CKKS/BFV encryption, as the random sample $w$ is unknown and the random sample $r$ is no longer accessible after the first modification. As a result, the distribution becomes independent of $\mu$, establishing semantic security.
[…]
@@
where $e_{r}=\langle\mathcal{H}(b^{\prime}-b),\mathbf{e}\rangle$, $\|e_{r}\|_{\infty}\leq\tau N\cdot B_{\chi}B_{\mathcal{H}}$. Combining these results, we can obtain that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ and $\langle sk,ct\rangle+\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle={\color[rgb]{1,0,0}\lfloor(Q/t)\cdot}\mu{\color[rgb]{1,0,0}\rceil}+e^{*}\approx{\color[rgb]{1,0,0}\lfloor(Q/t)\cdot}\mu{\color[rgb]{1,0,0}\rceil},$ where $\|e^{*}\|_{\infty}\leq(2N^{2}+4N)B_{\chi}+\tau N\cdot B_{\chi}B_{\mathcal{H}}$, as required.
V-B SMHE Construction
@@
    $C\leftarrow\texttt{SMHE.Encrypt}(\mu,pk)$: This algorithm encrypts the given plaintext $\mu\in R$ ($\mu\in R_{t}$) into the ciphertext $ct$ and generates masking ciphertexts $\{cx,\bm{\Gamma}\}$ under the encryption key $pk$, producing a tuple $C=\{ct,cx,\bm{\Gamma}\}$. The encryption and masking components are produced using the functions $\texttt{UniEnc}(\mu,pk)$ and $\texttt{MaskEnc}(r,pk)$, respectively. Specifically, the algorithm samples $w,r\leftarrow\chi$ and $e_{v0},e_{v1},e_{r0},e_{r1}\leftarrow D_{\sigma}$ and computes $ct=w\cdot pk+(\mu+e_{v0},e_{v1})\ (\text{mod }Q)$, $cz=r\cdot pk+(e_{r0},e_{r1})\ (\text{mod }Q)$ ($ct=w\cdot pk+(\lfloor(Q/t)\cdot\mu\rceil+e_{v0},e_{v1})\ (\text{mod }Q)$, $cz=r\cdot pk+(\lfloor(Q/t)\cdot\mu\rceil+e_{r_{0}},e_{r_{1}}))\ (\text{mod }Q)$), and $\bm{\Gamma}\leftarrow\texttt{GgtEnc}(sk,r)$.
@@
    $\overline{ct}:=\texttt{SMHE.Expand}(\{pk_{j}\}_{j\in[1,n]},i,ct)$: The algorithm expands the given ciphertext $ct=(c_{0},c_{1})$, encrypted under the public key $pk_{i}$, into an expanded ciphertext $\overline{ct}=(\bar{c}_{j})_{j\in[0,n]}$, where:
    $\bar{c}_{0}=c_{0}\quad\text{and}\quad\bar{c}_{j}=\begin{cases}c_{1}&\text{if }j=i,\\ 0&\text{otherwise},\end{cases}\quad\text{for}\ j\in[1,n].$
    We denote the associated key and the reference set of the expanded ciphertext $\overline{ct}$ as $pk_{i}$ and $\{pk_{j}\}_{j\in[1,n]}$, respectively.
@@
    $\overline{ct}_{\text{add}}\leftarrow\texttt{SMHE.Add}_{2}\big{(}\overline{ct}_{1},\overline{ct}_{2};pk_{1},pk_{2};\{cz_{i},$ $\bm{\Gamma}_{i}\}_{i\in[1,2]}\big{)}$: This algorithm takes as inputs two expanded ciphertexts $\overline{ct}_{1}=(c_{0}^{1},c_{1}^{1},0),\overline{ct}_{2}=(c_{0}^{2},0,c_{1}^{2})\in R_{q}^{3}$, which correspond to the fresh ciphertexts $ct_{1}=(c_{0}^{1},c_{1}^{1}),ct_{2}=(c_{0}^{2},c_{1}^{2})$, encrypted under keys $pk_{1}$ and $pk_{2}$, respectively. The associated masking parameters are $\{cz_{i},\bm{\Gamma}_{i}\}_{i\in[1,2]}$. The algorithm first performs $cx_{1}=(x_{0}^{1},x_{1}^{1})\leftarrow\texttt{Extend}(\bm{\Gamma}_{1},pk_{1},pk_{2}),cx_{2}=(x_{0}^{2},x_{1}^{2})\leftarrow\texttt{Extend}(\bm{\Gamma}_{2},pk_{2},pk_{1})$. Then, it masks the two ciphertexts as $\widehat{ct}_{1}=(c_{0}^{1}+x_{0}^{1}+z_{0}^{2},c_{1}^{1},z_{0}^{1}),\widehat{ct}_{2}=(c_{0}^{2}+x_{0}^{2}+z_{0}^{1},z_{0}^{2},c_{1}^{2})$111 The masked ciphertexts $\widehat{ct}_{1},\widehat{ct}_{2}$ can be equivalently interpreted as being obtained via an expansion of the masked fresh ciphertexts. That is, although the actual implementation performs masking after ciphertext expansion, the resulting ciphertexts $\widehat{ct}_{1},\widehat{ct}_{2}$ can be conceptually understood as the expansion of masked ciphertexts: $\widehat{ct}_{1}=\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,2]},1,ct_{1}+cx_{1}+cz_{2}),$ $\widehat{ct}_{2}=\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,2]},2,ct_{2}+cx_{2}+cz_{1}).$ This conceptual view clarifies the design intuition behind our masking strategy: masking is applied before expansion to hide sensitive plaintexts while enabling multi-key compatibility. . Finally, the algorithm outputs $\overline{ct}_{\text{add}}=\widehat{ct}_{1}+\widehat{ct}_{2}\pmod{Q}$.
@@
Algorithm 2 SMHE addition algorithm
1:$\overline{ct}=(c_{i})_{i\in[0,n]}$ and $\overline{ct}^{\prime}=(c^{\prime}_{i})_{i\in[0,n]}$, associated with public keys $\{pk_{j}\}_{j\in T}$ and $\{pk_{j}\}_{j\in T^{\prime}}$, respectively; The involved masking parameters $\{cz_{k},\bm{\Gamma}_{k}\}_{k\in T\cup T^{\prime}}$.
2:$\overline{ct}_{add}=(\bar{c}_{i})_{i\in[0,n]}\in R_{Q}^{n+1}$
3:$\bar{c}_{0}=c_{0}+c_{0}^{\prime}\pmod{Q}$
4:$\{cx_{i}=(0,0)\}_{i\in[1,n]}$, $\{cz_{i}^{\prime}=(0,0)\}_{i\in[1,n]}$
5:for $i\in[1,n]$ do
6:    if $i\in T$ then
9:    if $i\in T^{\prime}$ then
V-B1 Security of SMHE Construction

### v2

@@ 4.2 Security Vulnerability of CDKS
  • Decryption to obtain a global update: Each client $i$ first computes its partial decryption result $\nu_{i}\leftarrow\texttt{CDKS.PartD}$- $\texttt{ec}(\bar{c}_{i},s_{i})=c_{1}^{i}s_{i}+e_{i}\pmod{Q}(e_{i}\leftarrow D_{\sigma}),$ and broadcasts $\nu_{i}$ to the other clients or sends it to the aggregation server. Then, any client or the server computes the final decryption result as the global update:
    $\mu:=\texttt{CDKS.Merge}(\bar{c}_{0},\{\nu_{i}\}_{i\in[1,n]})={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\bar{c}_{0}+\sum_{i=1}^{n}\nu_{i}\big){\color[rgb]{1,0,0}\rceil}$
    $={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\sum_{i=1}^{n}(c_{0}^{i}+c_{1}^{i}\cdot s_{i})+\sum_{i=1}^{n}e_{i}\big){\color[rgb]{1,0,0}\rceil}\approx\sum_{i=1}^{n}\mu_{i}\ (\text{mod}\ Q)$
@@
Although the above process ensures correct parameter aggregation, it exposes local client parameters even under the honest-but-curious assumption, where both the server and the clients are assumed to faithfully execute the FL protocol while being curious to infer the privacy data of other clients. This security vulnerability stems from the ciphertext expansion and distributed decryption processes. Specifically, each component $\bar{c}_{i}$ ($i\in[1,n]$) of the expanded ciphertext $\overline{ct}=(\bar{c}_{0},\bar{c}_{1},\bar{c}_{2},\cdots,\bar{c}_{n})$ is associated with only the fresh ciphertext $c_{1}^{i}$ of the client $i$. This design facilitates decryption correctness, but introduces a critical security vulnerability. Since $c_{0}^{i}$ is publicly available – either directly from the fresh ciphertext $ct_{i}=(c_{0}^{i},c_{1}^{i})$ or from the expanded ciphertext $\overline{ct}_{i}$, and the partial decryption result $\nu_{i}$ is shared with the server or the clients for full decryption, any party with access to $\nu_{i}$ can trivially recover client $i$’s plaintext by computing ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(c_{0}^{i}+\nu_{i}\big){\color[rgb]{1,0,0}\rceil}$. That is,
${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(c_{0}^{i}+\nu_{i}\big){\color[rgb]{1,0,0}\rceil}={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(c_{0}^{i}+s_{i}c_{1}^{i}\big){\color[rgb]{1,0,0}\rceil}\approx\mu_{i}\pmod{Q}$, (3)
revealing the client $i$’s original plaintext $\mu_{i}$.
@@
5 New Multi-Key Variants of CKKS and BFV
To achieve our SMHE scheme, we retain the foundational structure of CDKS but introduce a masking scheme. This masking scheme ensures both the security and correctness of homomorphic addition, thereby enabling secure aggregation applications such as FL. Furthermore, ciphertexts processed through homomorphic addition with the masking scheme can still be directly evaluated by the homomorphic multiplication operations of CDKS without additional requirements or auxiliary computations.
5.1 A Masking Scheme for CKKS and BFV
We design a masking scheme tailored for CKKS and BFV, highlighting its role as a critical component in MKHE framework. Essentially, a masking scheme enables the use of a CKKS/BFV public key $pk$ (with an associated secret key $sk=(1,s)$) and a plaintext $\mu$ to generate a triple $(ct,cz,\bm{\Gamma})$. Here, $ct=(c_{0},c_{1})$ represents CKKS/BFV encryption of $\mu$ under $pk$, while $cz=(z_{0},z_{1})$ and $\bm{\Gamma}=(\bm{\varsigma}_{0},\bm{\varsigma}_{1})$ act as auxiliary information with two key properties: (1) the triple $(ct,cz,\bm{\Gamma})$ provides computational hiding for $\mu$, similar to $ct$ alone, and (2) when provided with another CKKS/BFV public key $pk^{\prime}$ (corresponding to a secret key $sk^{\prime}=(1,s^{\prime})$), it becomes feasible to construct a pair of polynomials $cx=(x_{0},x_{1})\in R_{Q}^{2}$ according to $\bm{\Gamma}$, such that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ and ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\langle sk,ct\rangle+\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\big){\color[rgb]{1,0,0}\rceil}\approx\mu\pmod{Q}$. The latter also implies that ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle(1,s,s^{\prime}),(c_{0}+x_{0}+z_{0},c_{1}+x_{1},z_{1})\rangle{\color[rgb]{1,0,0}\rceil}={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}\rangle{\color[rgb]{1,0,0}\rceil}\approx\mu\pmod{Q}$. Here, $\overline{sk}=(1,s,s^{\prime})$ is the concatenated secret key and $\overline{ct}=(c_{0}+x_{0}+z_{0},c_{1}+x_{1},z_{1})$ is a masked expanded ciphertext.
@@ 5.1 A Masking Scheme for CKKS and BFV
$(pp,pk,\texttt{UniEnc}(\mu,pk))$ $\approx_{c}(pp,pk,\texttt{UniEnc}(\mu^{\prime},pk)),$
$(pp,pk,\texttt{MaskEnc}(r,pk))$ $\approx_{c}(pp,pk,\texttt{MaskEnc}(r^{\prime},pk)),$
where $pp\leftarrow\texttt{FHE.Setup}(1^{\lambda})$, $(sk,pk)\leftarrow\texttt{FHE.Keygen}$ $(pp)$, $r^{\prime}\leftarrow\chi$, and $\mu^{\prime}\leftarrow R_{Q}$. FHE represents the traditional single-key CKKS/BFV scheme.
Correctness: Let $pp\leftarrow\texttt{FHE.Setup}(1^{\lambda})$, and consider two independently generated key pairs $(sk,pk)$ and $(sk^{\prime},pk^{\prime})$, obtained from $\texttt{FHE.Keygen}(pp)$. For any $\mu\in R_{Q}$, let $ct\leftarrow\texttt{UniEnc}(\mu,pk)$, $(cz,\bm{\Gamma})\leftarrow\texttt{MaskEnc}(r,pk)$, and $cx\leftarrow\texttt{Extend}(\bm{\Gamma},pk,pk^{\prime}).$ Then $\mu:=\texttt{FHE.Decrypt}(sk,ct)$ and
@@
    $ct$ $\leftarrow\texttt{FHE.Encrypt}(pk,\mu)$ (4)
    $=w\cdot pk+({\color[rgb]{1,0,0}\lfloor(Q/t)\cdot}\mu{\color[rgb]{1,0,0}\rceil}+e_{w_{0}},e_{w_{1}})\pmod{Q}$.
@@
      Sample $e_{r_{0}},e_{r_{1}}\leftarrow D_{\sigma}$ and encrypt the value 0 with $r$ and $pk$ using CKKS/BFV encryption algorithm and output
[…]
@@
The $\texttt{Extend}^{*}$ algorithm simplifies the process of running Extend for all $n$ public key $\{pk_{i}\}_{i\in[1,n]}$. Rather than executing $\texttt{Extend}(\bm{\Gamma},pk,pk_{1})+\cdots+\texttt{Extend}(\bm{\Gamma},pk,pk_{n})=\sum_{i\in[1,n]}(b_{i}-b)\boxdot\bm{\Gamma}$, it consolidates these computations into a single execution of $\texttt{Extend}^{*}(\bm{\Gamma},pk,\{pk_{i}\}_{i\in[1,n]})$.
Semantic Security. The attacker’s view consists of the distribution $\left(pp,pk,ct,cz,\bm{\Gamma}\right)$, where $pp\leftarrow\texttt{SMHE.Setup}(1^{\lambda})$, $(sk,pk)\leftarrow\texttt{SMHE.}\texttt{Keygen}(pp)$, $ct\leftarrow\texttt{UniEnc}(\mu,pk)$, $(cz,$ $\bm{\Gamma})\leftarrow\texttt{MaskEnc}(r,pk),\mu\in R_{Q}$. The semantic security proof leverages the security of CKKS/BFV encryption and follows these steps: (1) Modify $\bm{\Gamma}$ to the gadget encryption of random polynomial instead of the gadget encryption of $r$. (2) Replace $ct$ and $cz$ with encryptions of random messages. These are justified by the semantic security of CKKS/BFV encryption, as the random sample $w$ is unknown and the random sample $r$ is no longer accessible after encryption. Thus, the distribution becomes independent of $\mu$, establishing semantic security.
Correctness. Let $\{(sk,pk),(sk^{\prime},pk^{\prime})\}$ represent two valid key pairs produced by $\texttt{SMHE.Keygen}(pp)$. Recall that $sk=(1,s),sk^{\prime}=(1,s^{\prime});s,s^{\prime}\leftarrow\chi;pk=(b,a)\in R_{q}^{2},pk^{\prime}=(b^{\prime},a)\in R_{q}^{2}$ with $b=-a\cdot s+e\pmod{Q},b^{\prime}=-a\cdot s^{\prime}+e^{\prime}\pmod{Q}$, and $\|e\|_{\infty},\|e^{\prime}\|_{\infty}\leq\beta_{\chi}$.
@@
where $e_{r}=\langle\mathcal{H}(b^{\prime}-b),\mathbf{e}\rangle$ and $\|e_{r}\|_{\infty}\leq\tau N\cdot B_{\chi}B_{\mathcal{H}}$. Combining these results, we can obtain that $\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\approx 0$ and ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\langle sk,ct\rangle+\langle sk,cx\rangle+\langle sk^{\prime},cz\rangle\big){\color[rgb]{1,0,0}\rceil}={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}{\color[rgb]{1,0,0}\lfloor(Q/t)\cdot}\mu{\color[rgb]{1,0,0}\rceil}{\color[rgb]{1,0,0}\rceil}+e^{*}\approx\mu\pmod{Q}$, where $\|e^{*}\|_{\infty}\leq(2N^{2}+4N)B_{\chi}+\tau N\cdot B_{\chi}B_{\mathcal{H}}$, as required.
Algorithm 2 SMHE addition algorithm
[…]
5:for $i\in[1,n]$ do
6:  if $i\in T$ then
9:  if $i\in T^{\prime}$ then
5.2 SMHE Construction
@@
[…]
@@
    $\overline{ct}:=\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,n]},j,ct)$: The algorithm expands the given ciphertext $ct=(c_{0},c_{1})$, encrypted under $pk_{j}$, into an expanded ciphertext $\overline{ct}=(\bar{c}_{i})_{i\in[0,n]}$, where:
    $\bar{c}_{0}=c_{0}\quad\text{and}\quad\bar{c}_{i}=\begin{cases}c_{1}&\text{if }i=j,\\ 0&\text{otherwise},\end{cases}\quad\text{for}\ i\in[1,n].$
    We denote the associated key and the reference set of the expanded ciphertext $\overline{ct}$ as $pk_{j}$ and $\{pk_{i}\}_{i\in[1,n]}$, respectively.
@@
    $\overline{ct}_{\text{add}}\leftarrow\texttt{SMHE.Add}_{2}\big(\overline{ct}_{1},\overline{ct}_{2};pk_{1},pk_{2};\{cz_{i},$ $\bm{\Gamma}_{i}\}_{i\in[1,2]}\big)$: This algorithm takes as inputs two expanded ciphertexts $\overline{ct}_{1}=(c_{0}^{1},c_{1}^{1},0),\overline{ct}_{2}=(c_{0}^{2},0,c_{1}^{2})$ which are obtained by invoking $\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,2]},j,ct_{j})$ for $j\in[1,2]$. The associated masking parameters of the two parties are $\{cz_{i},\bm{\Gamma}_{i}\}_{i\in[1,2]}$. The algorithm first performs $cx_{1}=(x_{0}^{1},x_{1}^{1})\leftarrow\texttt{Extend}(\bm{\Gamma}_{1},pk_{1},pk_{2}),cx_{2}=(x_{0}^{2},x_{1}^{2})\leftarrow\texttt{Extend}(\bm{\Gamma}_{2},pk_{2},pk_{1})$. Then, it masks the two ciphertexts as $\widehat{ct}_{1}=(c_{0}^{1}+x_{0}^{1}+z_{0}^{2},c_{1}^{1},z_{0}^{1}),\widehat{ct}_{2}=(c_{0}^{2}+x_{0}^{2}+z_{0}^{1},z_{0}^{2},c_{1}^{2})$ [^2] The masked ciphertexts $\widehat{ct}_{1},\widehat{ct}_{2}$ can be equivalently interpreted as being obtained via an expansion of the masked fresh ciphertexts. That is, although the actual implementation performs masking after ciphertext expansion, $\widehat{ct}_{1}$ and $\widehat{ct}_{2}$ can be conceptually understood as the expansion of masked ciphertexts: $\widehat{ct}_{1}=\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,2]},1,ct_{1}+cx_{1}+cz_{2}),$ $\widehat{ct}_{2}=\texttt{SMHE.Expand}(\{pk_{i}\}_{i\in[1,2]},2,ct_{2}+cx_{2}+cz_{1}).$ This conceptual view clarifies the design intuition behind our masking strategy: masking is applied before expansion to hide plaintexts while enabling multi-key compatibility. . Finally, the algorithm outputs $\overline{ct}_{\text{add}}=\widehat{ct}_{1}+\widehat{ct}_{2}\pmod{Q}$.
@@
Remark: $\texttt{SMHE.Add}_{2}$ and SMHE.Add both denote the homomorphic addition of two given ciphertexts. The key difference lies in the underlying computation setting: $\texttt{SMHE.Add}_{2}$ corresponds to the two-party setting, where the reference set includes only the union of the two parties’ public keys; whereas SMHE.Add is defined in the general multi-party setting, where the reference set comprises the union of public keys from all multiple participating parties.
5.2.1 Security of SMHE